OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. sin 5x cos x - cos 5x sin x a.cos 6x b.cos 4x c.sin 6x d.sin 4x

OpenStudy (aravindg):

@hannahg what do you think ?

OpenStudy (anonymous):

I don't know how to start the problem :(

OpenStudy (aravindg):

do you know sin(A-B) formula ?

OpenStudy (anonymous):

no....

OpenStudy (anonymous):

sin (a-b)=sin(a) cos(b)- cos(a) sin(b)

OpenStudy (aravindg):

\[\sin(A-B)=\sin A \cos B -\cos A \sin B \]

OpenStudy (aravindg):

doesnt this look similiar to our question ?

OpenStudy (anonymous):

ohhh i see. i'll try to solve it:) thankyou!!

OpenStudy (anonymous):

@AravindG so i got sin(4x)

OpenStudy (aravindg):

very good!! well done ! :)